Find the maximum height of the curve
1
Expert's answer
2021-09-29T00:28:41-0400

Find the maximum height of the curve "y = 4\sin x - 3\cos x" above the "x" axis.


"\sqrt{(4)^2+(3)^2}=5"

"y =5( \dfrac{4}{5}\sin x - \dfrac{3}{5}\cos x)"

Let "\cos \theta=\dfrac{4}{5}, \sin\theta=\dfrac{3}{5}, \theta=\sin^{-1}(\dfrac{3}{5})."

Then

"\dfrac{4}{5}\sin x - \dfrac{3}{5}\cos x=\sin(x-\theta), \theta=\sin^{-1}(\dfrac{3}{5})"

"y=5\sin^{-1}(\dfrac{3}{5})"

"-1\leq\sin^{-1}(\dfrac{3}{5})\leq 1"

"-5\leq y\leq 5"

The maximum height of the curve of the curve "y = 4\sin x - 3\cos x" above the "x" axis is "5."